2 cyclists leave one is traveling east and the other west. after 4 hours they are 112 miles apart. if one cyclist is traveling 4mph faster than the other find the speed of each cyclist.
Answer by rfer(16322)

(x+x+4)(4)=112
(2x+4)(4)=112
8x+16=112
8x=96
x=12 mph
x+4=16 mph